OpenStudy (anonymous):

Simplify... (x^2-x-30)/(x^2+13x+40)*(x^2+11x+24)/(x^2-9x+18) (The division symbols show fractions.)

OpenStudy (anonymous):

\[\frac{x^2-x-30}{x^2+13x+40}\cdot\frac{x^2+11x+24}{x^2-9x+18}\] Factor: \[\frac{(x-6)(x+5)}{(x+8)(x+5)}\cdot\frac{(x+8)(x+3)}{(x-6)(x-3)}\] Cancel out any terms that can be cancelled.

OpenStudy (anonymous):

(x+3)/(x-3)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank You!